A bar magnet makes 40 oscillations per minute in an oscillation magnetometer. An identical magnet is demagnetized completely and is placed over the magnet in the magnetometer. Find the time taken for 40 oscillations by this combination. Neglect any induced magnetism.

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. γ base 1 = 40 oscillations/minute B base H = 25 μT m fo second magnet = 1.6 A-m^2 d = 20 cm = 0.2 m (a) For north facing north γ base 1 = 1/2π√MB base H/I γ base 2 = 1/2π√M(B base H – B)/I B = μ base 0/4π m/d^3 = 10^-7 *1.6/8 * 10^-3 = 20 μT γ base 1/γ base 2 = √B/B base H – B ⇒ 40/γ base 2 √25/5 ⇒ γ base 2 = 40/√5 = 17.88 = 18 osci/min (b) For north pole facing south γ base 1 = 1/2π√MB base H/I γ base 2 = 1/2π√M(B base H – B)/I γ base 1/γ base 2 = √B/B base H – B ⇒ 40/γ base 2 = √25/45 ⇒ γ base 2 = 40/√(25/45) = 53.66 = 54 osci/min